In adding or subtracting fractions ANY common denominator will allow you to add /subtract the two fractions. However discuss what is the advantage might be in using the LEAST common denominator to add/subtract them.
Answer by Alan3354(69443)

There is no advantage.
--------
eg, 1/6 + 1/4
Use the produce 6*4 = 24
--> 4/24 + 6/24 = 10/24 = 5/12
============
Zero time spent finding the LCD.
Reducing from 10/24 to 5/12 is easy.
===========
IDK why people fret about the LCD.